# Universal Approximation Theory: The basic theory for deep learning-based computer vision models

### v2

@@ 4.1 Why must CNNs be Deep?
The second reason is derived from the UAT. As shown in Figure 1, increasing the number of layers in the network means a larger $N$, which brings the UAT closer to approximating the target function. An image can be understood as a special function in a high-dimensional space, exhibiting strong correlations in 2D or 3D space. These two reasons are also applicable to ViTs. Figure 7 shows the preprocessing of ViTs, where $\tilde{\mathbf{x}}$ is the input to the ViTs. It is evident that each row shares the same parameters, and the learning of each row adheres to the principles of local receptive fields and spatial invariance.
@@ 4.2 Why Residual-Based CNNs Excel in CV: Superior Generalization Ability
Figure 8: VGG Vs ResNet: The characters in ’()’ represent the corresponding operation’s parameters. Hereinafter the same.
In the previous discussion, we explained why deep networks are essential for learning in CNNs. However, why did VGG, despite being a deep network, not dominate the subsequent development of CV, whereas residual-based CNNs did? ResNet significantly shaped the design of later network architectures, with the residual structure appearing in many subsequent networks. What exactly endows the residual structure with such powerful capabilities? We think the answer is the generalization capability given by residual structure. To analyze the source of this power, we use the Matrix-Vector Method to express VGG and the residual structure in equations, and then compare them with the UAT.
Figure 8 illustrates the basic structures of VGG and ResNet. Based on this figure, a $i$-layer VGG can be written as:
@@
which perfectly aligns with the multilayer UAT mathematical form proposed by Hornik (Hornik, Stinchcombe, and White 1989). On the other hand, a multi-layer residual-based CNN can be expressed as:
where $\mathbf{b}^{\prime}_{j,1}=(\mathbf{W}^{\prime}_{j,1}\mathbf{b}^{\prime}_{j-1,2}+\mathbf{b}^{\prime}_{j,1})+\mathbf{W}^{\prime}_{j,1}UAT_{j-1}^{R}$ and $UAT_{j-1}^{R}=\Sigma_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,2}\sigma(\mathbf{W}^{\prime}_{k,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,1})$ (see Appendix A.2 for detail). Eq. (4) conforming to the UAT form given in Figure 1. In this equation, $\mathbf{b}_{j,1}$ is a dynamic parameter approximated by $j$ layers UAT ($j>1$). It is evident that both VGG and ResNet are concrete implementations of UAT. Theoretically, their approximation capabilities should be equivalent, and any differences could be mitigated by increasing the number of network layers. So why is ResNet significantly more powerful than VGG?
The primary reason lies in the dynamic approximation capabilities of residual networks. As seen in Eq. (3), once a VGG network is trained, its corresponding UAT parameters are fixed. This means that a VGG network can only approximate a single, fixed function. Although UAT has strong approximation abilities, image data is highly variable, and the corresponding function often changes. In contrast, residual networks can dynamically approximate the corresponding function based on the input because the biases in residual networks are influenced by the input.
Therefore, while the original UAT formula has strong approximation capabilities, using it for approximation (equivalent to using a VGG network in CV) means that once training is complete, the UAT parameters are fixed, and the function it can approximate is also fixed. In contrast, in residual networks, the UAT parameters can change based on the input, allowing the function being approximated to vary accordingly. This adaptability is the fundamental source of the superiority of residual networks.
@@ 4.3 The Difference between Residual-Based CNNs and ViTs
Figure 9: The process of Transformer.
In this section, we will compare the UAT format of Transformer-based ViTs and residual-based CNNs. Figure 9 shows the general form of a Transformer. A multi-layer Transformer can be expressed as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (5)
where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$ and $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$ (see Appendix A.3 for details). The mathematical representation of ViT also conforms to UAT, with $\mathbf{b}^{\prime}_{j,2}(j>1)$ is approximated by $j$-layer UAT. The biggest difference between Transformers and residual-based CNNs lies in the parameters in the MHA mechanism dynamically changing with the input. Therefore, in the corresponding UAT mathematical form for Transformers, all $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ parameters for the $i$-th layer, where $j=1,...,i$, are dynamically changing with the input.
@@
The main difference between Transformer networks and residual-based multi-layer convolutional networks lies in how the input affects their corresponding UAT parameters. In Transformer networks, input information influences both weight and bias in the corresponding UAT. In contrast, in networks constructed with multiple residual convolutions, input information only influences bias in the corresponding UAT. Essentially, both types of networks have the ability to dynamically approximate functions based on input. This aligns with their nearly identical performance in the field of CV, supporting the theoretical foundation that both are specific implementations of UAT capable of dynamically approximating the corresponding functions based on input.
4.4 Human Vision vs UAT Vision
@@
Based on the above discussion, we can consider that deep learning-based CV models are implementations of UAT, which we refer to as UAT vision. So, what is the relationship between human vision and UAT vision? We believe that the human brain functions as a large network composed of one or multiple UATs. Therefore, the concepts of memory and understanding, as defined by humans, do not fundamentally exist; rather, they can be categorized as reasoning based on the UAT in our brains. Just as deep vision networks do not have a database storing images, they directly approximate the corresponding results from the inputs (which can be simplified as $y=UAT(x)$). Specifically, humans and animals are born with certain weights in their brains, which are innate reflexes, such as infants instinctively knowing how to suckle. However, other weights in the brain need to be learned, and our process of observing the world from a young age is essentially a process of training these weights. Given that the eyes’ vision operates at 60 frames per second, the brain would be trained on hundreds of millions of high-resolution images over a year.
@@
5 Conclusion
This paper delves into the theoretical foundations of deep learning in the field of CV. Specifically, the current CV landscape is primarily dominated by CNNs and Transformer models. We utilize the Matrix-Vector Method to unify these models under the framework of the UAT. Based on this framework, we provide explanations for several common issues and techniques in CV. The requirement for deep networks in CNNs is jointly determined by the intrinsic characteristics of image data and the demands of UAT theory. The robust generalization ability of residual networks stems from their ingenious design, which enables UAT to dynamically adapt to corresponding functions based on input data. Similarly, Transformer-based models possess this ability, with the key difference lying in which parameters of UAT they influence. Residual networks primarily affect the bias term, while Transformers influence both the weights and the bias. Additionally, we think the thinking process of humans in images is the same as UAT, they all use pre-trained parameters to fit the output. But the human brain is a more complex UAT system and it is on training all the time.
@@ Appendix A The UAT Format of Residual-based CNNs and Transformer-based ViTs
Note: The sections A.1 and A.3 have already been proven in UAT2LLMs. For completeness, we also include them in this paper.
In this section, we will use the matrix-vector method to demonstrate the mathematical forms of residual-based CNNs and Transformer-based ViTs, and their relationship with UAT. In Section A.1, we first presented some lemmas of UAT. Using these lemmas, we then proved that residual-based CNNs A.2 and Transformer-based ViTs A.3 are specific implementations of UAT.
@@ A.1 The Properties of UAT
Before expressing CNNs based on residuals and Transformers based on ViTs in the UAT format, we present two lemma regarding UAT. There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$. The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$. Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
@@
Equation 1 shows the representation of UAT multiplying a matrix. Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the result remains consistent with the original UAT mathematical form. Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
Lemma 2. Let the general term of the network be written as: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, $i=0,1,2,3,\ldots$. The multi-layer network of this form still corresponds to the mathematical form of UAT, and we refer to such a network as a residual network.
To prove that a multi-layer network with the general term written as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ corresponds to the UAT mathematical form, we first provide the forms of single-layer and two-layer networks, as shown in Equations (2) and (3).
@@
In Equation (3), let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Equation (4) can be written as:
@@
Why can we combine the above terms? Because once the network training is complete, we know these parameters, so they can be directly calculated and thus combined. And $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$ clearly has the same mathematical form as UAT. We can understand it as dynamically fitting the bias term $\mathbf{b}^{\prime}_{2,2}$ using UAT based on the input. Therefore, we have proven that the mathematical form of one and two-layer residual networks is consistent with the UAT mathematical form.
@@
Assume that the mathematical form of the first $i$ layers of the residual network is consistent with UAT. Our goal is to prove that the mathematical form of the $i+1$-th layer of the residual network is still consistent with UAT. For convenience of expression, we make the following definition: since the mathematical form of the first $i$ layers of the residual network is consistent with UAT, we write $\mathbf{x}^{\prime}_{i}$ as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, where the first term is written separately and the rest is written as the remainder term $UAT^{R}_{i}$. Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
@@
First, consider the part $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$. Substitute $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$ into it, and we get Equation 5. Let $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with UAT, and we have already proven that the mathematical form of UAT remains unchanged when multiplied by a matrix, we have proven that the mathematical form of the first part is consistent with UAT.
@@
Next, we prove that the mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is consistent with UAT. By substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$ into it, we obtain Equation 6. Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$. Thus, the second term can be written as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be considered as a term within UAT. Here, $\mathbf{b}^{\prime}_{i+1,2}$ is a bias term approximated using UAT.

### v4

@@ 4.1 Why must CNNs be Deep?
The second reason is derived from the DUAT. As shown in Figure 1, increasing the number of layers in the network means a larger $N$, which brings the DUAT closer to approximating the target function. An image can be understood as a special function in a high-dimensional space, exhibiting strong correlations in 2D or 3D space. These two reasons are also applicable to ViTs. Figure 8 shows the preprocessing of ViTs, where $\tilde{\mathbf{x}}$ is the input to the ViTs. It is evident that each row shares the same parameters, and the learning of each row adheres to the principles of local receptive fields and spatial invariance.
@@ 4.2 Why Residual-Based CNNs Excel in CV: Superior Generalization Ability
In the previous discussion, we explained why deep networks are essential for learning in CNNs. However, why did VGG, despite being a deep network, not dominate the subsequent development of CV, whereas residual-based CNNs did? ResNet significantly shaped the design of later network architectures, with the residual structure appearing in many subsequent networks. What exactly endows the residual structure with such powerful capabilities? We think the answer is the generalization capability given by residual structure. To analyze the source of this power, we use the Matrix-Vector Method to express VGG and the residual-based CNNs into equations, and then compare them.
Figure 10: VGG: The characters in ’()’ represent the corresponding operation’s parameters.
Figure 10 illustrates a basic structure of VGG. Based on this figure, a $i$-layer VGG can be written as:
@@
which perfectly aligns with the mathematical formulation of the multilayer UAT proposed by Hornik Hornik et al. (1989). Therefore, based on Eqs. 6 and 3, we can conclude that the fundamental reason ResNet is more powerful than VGG lies in the fact that the residual network is a DUAT function, while VGG is a UAT function. DUAT has a stronger dynamic approximation capability. As shown in Eq. 6, once the VGG network finishes the training, its corresponding UAT parameters become fixed. This means that the VGG network can only approximate a single, fixed function. Although UAT has strong approximation capabilities, image data is highly variable, and the corresponding functions often change.
In contrast, the residual network, corresponding to the DUAT in Eq. 3, can dynamically adjust its bias parameters based on the input, allowing it to approximate the corresponding function dynamically. This adaptability is the fundamental source of the superiority of residual networks.
@@ 4.3 The Difference between Residual-Based CNNs and ViTs
In this section, we will compare the DUAT format of Transformer-based ViTs and residual-based CNNs. Figure 9 shows the general form of a Transformer. A multi-layer Transformer is shown in Eq 4. Comparing Eq. 4 and 3, the main difference between Transformer networks and residual-based multi-layer convolutional networks lies in how the input affects their corresponding DUAT parameters. In Transformer networks, input information influences both weight and bias in the corresponding DUAT. In contrast, in networks constructed with residual-based CNNs, input information only influences bias in the corresponding DUAT. Essentially, both types of networks have the ability to dynamically approximate functions based on input. This aligns with their nearly identical performance in the field of CV, supporting the theoretical foundation that both are DUAT functions and capable of dynamically approximating the corresponding functions based on input.
5 Conclusion
This paper delves into the theoretical foundations of deep learning in the field of CV. Specifically, the current CV landscape is primarily dominated by CNNs and Transformer models. We utilize the Matrix-Vector Method to unify these models under the framework of the DUAT. Based on this framework, we provide explanations for several common issues and techniques in CV. The requirement for deep networks in CNNs is jointly determined by the intrinsic characteristics of image data and the demands of DUAT. The robust generalization ability of residual networks stems from their ingenious design, which enables UAT to dynamically adapt to corresponding functions based on input data. Similarly, Transformer-based models possess this ability, with the key difference lying in which parameters of UAT they influence. Residual networks primarily affect the bias term, while Transformers influence both the weights and the bias.
6 Some Guesses on Human Vision Based on DUAT
@@
Based on the above discussion, we can consider that residual-based deep learning CV models are DUAT functions, which we refer to as DUAT vision. So, what is the relationship between human vision and DUAT vision? We believe that the human brain functions as a large network composed of one or multiple DUATs. Because first neural network is perceptron Rosenblatt (1963) which was designed based on the brain neuron. The Transformer and residual-based CNNs are DUAT functions, and the perceptron can be seen as a special form of UAT. Therefore, the concepts of memory and understanding, as defined by humans, can be categorized as reasoning based on the DUAT in our brains. Just as deep vision networks do not have a database storing images, they directly approximate the corresponding results from the inputs (which can be simplified as $y=DUAT(x)$). Specifically, humans and animals are born with certain weights in their brains, which are innate reflexes, such as infants instinctively knowing how to suckle. However, other weights in the brain need to be learned, and our process of observing the world from a young age is essentially a process of training these weights. Given that the eyes’ vision operates at 30 frames per second Lu et al. (2017), the brain would be trained on hundreds of millions of high-resolution images over a year.
@@ Appendix A The UAT Format of Residual-based CNNs and Transformer-based ViTs
Note: The sections A.1, Section A.2 and A.4 have already been proven in DUAT2LLMs. We also include them in this paper to ensure completeness.
In this section, we will show the DUAT mathematical form corresponding to residual-based CNNs and Transformer-based ViTs (which also use residual connections by design). In Section A.1, we first introduce a lemma concerning UAT and then we introduce the DUAT in Section A.2. Based on Section A.2, we show that both residual-based CNNs (see Section A.3) and Transformer-based ViTs (see Section A.4) are DUAT functions.
It is important to highlight that DUAT’s overall mathematical structure is aligned with UAT; the key distinction is that DUAT’s parameters are influenced by the input. Currently, this influence is achieved primarily through the use of residual connections. Therefore, we represent multi-layer residual networks in the corresponding UAT form and illustrate how their parameters are influenced by the input, establishing them as DUAT functions. Due to the close relationship between UAT and DUAT, some interchange of the terms UAT and DUAT may occur in the following sections.
@@ A.1 The Properties of UAT
Before expressing residual-based CNNs and Transformer-based ViTs in the UAT format, we present a lemma regarding UAT. There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$. The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$. Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
@@
Eq. 7 shows the representation of UAT multiplying a matrix. Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the general mathematical form of Eq. 7 remains consistent with the original UAT mathematical form. Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
A.2 From UAT to DUAT
In this section, we first present a general form for a single-layer residual term in a network and demonstrate that the mathematical form of a multi-layer network composed of this residual term aligns with the DUAT framework. Before proceeding with the formal proof, we first define DUAT explicitly: DUAT shares the overall mathematical form of UAT, but with certain parameters influenced by the input, allowing them to change dynamically in response to it. We refer to these parameters as ”dynamic parameters.” In standard UAT, parameters remain fixed once training is complete, whereas, in DUAT, the dynamic parameters are functions of the input and thus vary with it. Consequently, dynamic parameters in DUAT may take the form of complex functions. Currently, DUAT primarily is implemented by residual structure, meaning that, in general, these complex functions are also DUAT. Next, we will give the proof.
A general residual term of the network can be written as:
$\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,2})$ (8)
where $i=1,2,3,\ldots$ and $\mathbf{x}^{\prime}_{0}$ represents the input. A multi-layer network structured in this way aligns with the mathematical form of DUAT. To demonstrate that a multi-layer network leverages the general term of Eq. 8 corresponds to DUAT, we first examine the forms of single-layer and two-layer networks, as shown in Eqs. 9 and 10.
@@ A.2 From UAT to DUAT
In Eq. 10, let $\mathbf{W}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{2,2}=\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$. Thus, Eq. 10 can be written into Eq. 11.
@@
For the cleatity, we could write Eq. 9 and Eq. 11 into Eq. 12 and Eq. 13, where $UAT^{R}_{1}=\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ and $UAT^{R}_{2}=\Sigma_{j=1}^{2}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$.
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+UAT^{R}_{1}$ (12)
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+UAT^{R}_{2}$ (13)
According to Eq. 12 and Eq. 13, aside from the initial term, the overall mathematical forms of the residual terms $UAT^{R}_{1}$ and $UAT^{R}_{2}$ are consistent with UAT. However, for $UAT^{R}_{2}$, its parameter $\mathbf{b}^{\prime}_{2,2}$ is influenced by the input, while other weight parameters, such as $\mathbf{W}^{\prime}_{2,1}$, can be disregarded. These parameters are only affected by other parameters, so once the network training is complete, they are essentially fixed. This allows us to focus primarily on the dynamic parameters. In Eq. $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, except for the initial term, $\mathbf{b}^{\prime}_{2,2}$ also follows the same mathematical form as UAT. This can be interpreted as dynamically adjusting the bias term $\mathbf{b}^{\prime}_{2,2}$ via UAT based on the input. In conclusion, we have demonstrated that the mathematical forms of one-layer and two-layer residual networks are consistent with the UAT framework and the two-layer residual network is the DUAT function.
Next, we will use mathematical induction to prove that the mathematical form of a multi-layer residual network is also a DUAT function. Assume that the overall mathematical form of the first $i$ layers of the residual network aligns with UAT. Our goal is to show that the overall mathematical form of the $i+1$-th layer remains consistent with UAT.
@@
For clarity, we make the following definitions: since the overall mathematical form of the first $i$ layers is consistent with UAT, we can write it as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, where the first term is explicitly written, and the remainder is denoted as $UAT^{R}_{i}$, with $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$. Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
@@
First, consider $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$. Substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, we obtain Eq. 14. Setting $\mathbf{W}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}$ and $\mathbf{b}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$. Since the overall mathematical form of $UAT^{R}_{i}$ is consistent with the $i$-layer UAT and we have shown that the UAT form is preserved when multiplied by a matrix, we have thus demonstrated that the overall mathematical form of the first part remains consistent with UAT.
@@
Next, we show that the overall mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is also consistent with UAT. Substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$ into this expression, we arrive at Eq. 15. Setting $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}$ and $\mathbf{b}^{\prime}_{i+1,2}=(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}$, we can rewrite the second part as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be interpreted as a term in UAT. Here, $\mathbf{b}^{\prime}_{i+1,2}$ serves as a bias term approximated using DUAT.